OpenStudy (sharkfinsoup):

A boy slides a book across the floor, using a force of 5 N over a distance of 2 m. What is the kinetic energy of the book after he slides it? Assume there is no friction. I have no idea what formula to use. Can someone point me in the right direction?

OpenStudy (osprey):

work is force times distance. To give something energy, work has to be done. the boy gives 5N of force effort over 2m. That's 10 joules of mech effort. no friction means that all the work goes into increasing the kinetic energy. So the ke is 10 joules. If the mass were given in the q, then the actual speed could be calculated from 1/2 m v^2. Seems as though this site is closing end of jan, so I've been flashed. http://perendis.webs.com

OpenStudy (sharkfinsoup):

Thank you so much. I really appreciate it

OpenStudy (osprey):

you're more than welcome

OpenStudy (osprey):

@sharkfinsoup = shark fin soup ?

OpenStudy (sharkfinsoup):

yeah I just made it randomly @osprey

OpenStudy (osprey):

I quite like the sound of the soup ... Fish soup is, I believe, a french delicacy. Were you in a posh restaurant, poss in Caribbean at the time, or were you near some financial speculators and other sales persons ? my own is almost the same ... 'cepting it's actually the name of the wireless device I use.

OpenStudy (sharkfinsoup):

No I just made this account today and thats the first thing that popped in my head @osprey

OpenStudy (vipulrakta):

SIMPLY APPLY WORK ENERGY THEREM WORK DONE = CHANGE IN K.E . \[W= \Delta K.E \] \[W=1/2M(V _{1}^{2}-V _{2^{2}})\] \[V _{2}=0\] \[1/2(MV ^{2} ) = W = F*D =5N*2M = 10J\]